### v1

# Exponential Dichotomies in Higher-Dimensional Spatial Dynamics for Elliptic Partial Differential Equations

Abstract
Exponential dichotomies, when they exist, provide powerful information about the structure of bounded solutions even in the case of an ill-posed evolutionary equation. The method of spatial dynamics, in which one views a spatial variable as a time-like evolutionary variable, allows for the use of classical dynamical systems techniques, such as exponential dichotomies, in broader contexts. This has been utilized to study stationary, traveling wave, time-periodic, and spiral wave solutions of PDEs on spatial domains with a distinguished unbounded direction (e.g. the real line or a channel of the form $\mathbb{R}\times\Omega$). Recent work has shown how to extend the spatial dynamics framework to elliptic PDEs posed on general multi-dimensional spatial domains. In this paper, we show that, in the same context, exponential dichotomies exist, thus allowing for their use in future analyses of coherent structures, such as spatial patterns in reaction-diffusion equations on more general domains.
Acknowledgements: This material is based upon work supported by the National Science Foundation under Award Numbers DMS-2205434, DMS-2006887, and DMS-2307650.
1 Introduction
A current challenge in the study of pattern formation is the intractability of analytic dynamical systems approaches for pattern forming partial differential equations in higher spatial dimensions. Dynamical systems techniques, for example center manifold reductions, geometric singular perturbation theory for fast-slow systems, and exponential dichotomies, have been applied to study existence and stability of stationary spatial patterns and traveling wave solutions in a variety of one-dimensional contexts (e.g. [4, 13, 16]).
Spatial dynamics, in its simplest iteration, is the trick of rewriting a second order differential equation into a system of first order equations. For example, $u_{xx}+F(u)=0$ becomes the system
$\begin{array}[]{l}u_{x}=v\\ v_{x}=-F(u).\end{array}$
For PDEs posed on $\mathbb{R}$, this perspective allows stationary and traveling-wave problems to be studied as an ODE, permitting the use of well-understood and widely applicable dynamical system tools. Similarly, PDEs posed on $\mathbb{R}^{n}$ that are restricted to radially symmetric solutions depend only on the radius $r\in\mathbb{R}^{+}$ and can therefore be written as an ODE using spatial dynamics. For example, $\Delta u+F(u)=0$, when written in spherical coordinates with $u=u(r)$, becomes
$\quad u_{rr}+\left(\frac{n-1}{r}\right)u_{r}+F(u)=0$
which is then written as a first order system in the variable $r$ as
$\begin{array}[]{l}u_{r}=v\\ v_{r}=-\left(\frac{n-1}{r}\right)v-F(u).\end{array}$
The nonautonomous $r$-dependence of this radial equation and the singularity that occurs at $r=0$ make this setting meaningfully different from the one-dimensional framework. Scheel made significant contributions to characterizing the behavior of radially symmetric spatial patterns in [19] where he developed a systematic bifurcation theory including extensions of center manifold reduction and normal form theory. In [19], Scheel also established existence of both stationary and time-periodic radially symmetric patterns using these techniques.
When spatial dynamics is applied to elliptic a PDEs posed on a cylindrical domain $\mathbb{R}\times\Omega\subset\mathbb{R}^{n}$, the result is an evolutionary equation posed on an infinite-dimensional function space (rather than a finite dimensional vector space as in the case of ODEs). The same reaction diffusion problem $\Delta u+F(u)=0$ is decomposed into the variables $(x,y)\in\mathbb{R}\times\Omega$ to become
$u_{xx}+\Delta_{\Omega}u+F(u)=0,\qquad u(x,\cdot)\in L^{2}(\Omega).$
The spatial evolutionary equation that results from writing the above as a first-order system in $x$ is
$\begin{array}[]{l}u_{x}=v\\ v_{x}=-\Delta_{\Omega}u-F(u).\end{array}$
There is a long history of extending dynamical systems tools into this infinite-dimensional context (e.g. [6, 7, 8, 11]). These extensions have made possible existence proofs in the channel setting [5], construction of local and global manifolds [14], alternative methods to understand the Fredholm properties of elliptic operators [15], and stability analyses of traveling waves [13].
A similarly infinite-dimensional evolutionary equation emerges when a spatial dynamics perspective is applied to problems posed on $\mathbb{R}^{2}$, written in spherical coordinates $(r,\varphi)\in\mathbb{R}^{+}\times S^{n-1}$. This was first explored in the study of spiral wave patterns by Scheel in [18]. Later work by Sandstede and Scheel treated spiral waves in greater detail. They applied a dynamical systems perspective, and in particular the existence of exponential dichotomies, to rigorously characterize the spectral properties of spiral waves on both bounded and unbounded domains, the robustness of spiral waves, and the behavior of eigenfunctions to the linearization which indicate the spatial structure of patterns bifurcating from radially symmetric spiral waves [17].
Recent work by Beck et. al. [2] has extended the use of spatial dynamics further to arbitrary Euclidean spatial domains, and in particular to elliptic PDEs posed on $\mathbb{R}^{n}$. The application of spatial dynamics in $\mathbb{R}^{n}$ looks much the same as in the setting of $\mathbb{R}^{2}$, relying again on spherical coordinates $(r,\varphi)\in\mathbb{R}^{+}\times S^{n-1}$. The elliptic PDE
$0=\Delta u-V(x)u,\qquad V:\mathbb{R}^{n}\to\mathbb{C}^{d\times d},\quad u\in\mathbb{R}^{n}\to\mathbb{C}^{d}$ (1)
is rewritten in spherical coordinates as
$u_{rr}+\left(\frac{n-1}{r}\right)u_{r}+\frac{1}{r^{2}}\Delta_{S^{n-1}}u+V(r,\varphi)u=0,\qquad u(r,\cdot)\in L^{2}(S^{n-1},\mathbb{C}^{d})$
where $\Delta_{S^{n-1}}$ is the spherical Laplacian (well understood via spherical harmonics). The above can be expressed as a first order system in $r$ as
$\mathbf{u}_{r}=\mathcal{A}(r)\mathbf{u}=\begin{pmatrix}0&1\\ V(r,\varphi)-\frac{1}{r^{2}}\Delta_{S^{n-1}}&-\frac{(n-1)}{r}\end{pmatrix}\mathbf{u}$ (2)
posed on the Banach space $X=H^{1}(S^{n-1},\mathbb{C}^{d})\times L^{2}(S^{n-1},\mathbb{C}^{d})$ so that $\mathbf{u}=(u,v)\in X$.
The results of Beck et. al. establish not only the method of spatial dynamics, but also the equivalence of weak PDE solutions (e.g. weak solutions to (1)) and solutions to the corresponding spatial evolutionary equation (e.g. solutions to (2)) in both $\mathbb{R}^{n}$ and general bounded Euclidean domains. These results lay the groundwork for adapting classical dynamical systems techniques to this newer spatial dynamics setting. In fact, in [1], Beck et. al. show the existence of an exponential dichotomy for $r\in(0,R]$ (the core field) and prove that the unstable space defined by this dichotomy corresponds to the boundary data of weak PDE solutions; there is a correspondence between “well-behaved” solutions of the spatial evolutionary system and weak PDE solutions.
Exponential dichotomies extend the notion of invariant stable and unstable linear subspaces to the nonautonomous and infinite dimensional settings [3, 9, 10, 11]. They characterize a hyperbolic splitting into an exponentially, asymptotically stable invariant evolution and an exponentially, asymptotically unstable invariant evolution. The incredible power of exponential dichotomies is that they allow us to characterize asymptotic behavior in systems where other tools for asysmptotic analysis may fail. For example, even for spatial evolutionary equations that are ill-posed, exponential dichotomies may exist and can then describe the behavior of bounded solutions. Exponential dichotomies have been used to understand spectral and Fredholm properties of operators (e.g. [3, 10, 15, 17]), to construct center manifold reductions (e.g. [6, 7, 8, 17, 18]), and to construct stable and unstable manifolds and therefore to also study homoclinic and heteroclinic solutions (e.g. [9, 12, 16]).
In this paper, we prove the existence of a far field dichotomy— an exponential dichotomy on the interval $r\in[R,\infty)$ for $R<\infty$— for the same spatial evolutionary equation (2). Existence proofs for exponential dichotomies typically rely heavily on the robustness of exponential dichotomies which was first established in an infinite dimensional setting by Peterhof, Sandstede, and Scheel in [11]. This property means that exponential dichotomies persist through small perturbation, so existence arguments are often reduced to finding the “right” decomposition of an operator into a principal part that is known to have an exponential dichotomy and a perturbation that is sufficiently small.
In [1], the primary challenge of proving the existence of the core-field dichotomy lay in the singularity at $r=0$ which was overcome by a reparameterization of $\tau=\ln r$ (see also [18]) and an exponential rescaling of the resulting system posed on $\tau\in\mathbb{R}^{-}$. In this new rescaled system, the asymptotic operator that emerges when $\tau$ is sent to -$\infty$ can be chosen to be the principal operator while the perturbation that remains decays exponentially in $\tau$.
Applying robustness to prove the existence of a far-field exponential dichotomy is quite challenging due to the nonuniform unboundedness of the $r^{-2}\Delta_{S^{n-1}}$ term. Though this term vanishes as $r$ is sent to infinity and does not primarily control the hyperbolicity of the system, its unboundedness with respect to the $L^{2}$ norm means that it cannot simply be considered a “perturbation.” Instead, we must consider this Laplacian inside of the principal part and work within a function space where it is bounded but does not dominate the underlying hyperbolicity of the system. Because of the nonuniformity of the $r^{-2}\Delta_{S^{n-1}}$ term, we equip the underlying function space $X=H^{1}(S^{n-1})\times L^{2}(S^{n-1})$, on which (2) is posed, with an explicitly $r$-dependent norm, taking inspiration from the approach of Sandstede and Scheel in the spiral waves context [18, 17].
Observing that $\mathcal{A}(r)$ is both closed and densely defined in $X$ with domain $X^{1}=H^{2}(S^{n-1})\times H^{1}(S^{n-1})$, we adapt the natural norms for $X$ and $X^{1}$ which are given by
$\|\mathbf{u}(r,\cdot)\|_{X}^{2}$ $=\|u(r,\cdot)\|_{H^{1}}^{2}+\|v(r,\cdot)\|_{L^{2}}^{2}$
$\|\mathbf{u}(r,\cdot)\|_{X^{1}}^{2}$ $=\|u(r,\cdot)\|_{H^{2}}^{2}+\|v(r,\cdot)\|_{H^{1}}^{2},$ (3)
to $r$-dependent norms by weighting the $H^{1}$ and $H^{2}$ norms, respectively:
$\|\mathbf{u}(r,\cdot)\|_{X_{r}}^{2}$ $:=\frac{1}{r^{2}}\|u(r,\cdot)\|_{H^{1}}^{2}+\|u(r,\cdot)\|_{L^{2}}^{2}+\|v(r,\cdot)\|_{L^{2}}^{2}$
$\|\mathbf{u}(r,\cdot)\|_{X^{1}_{r}}^{2}$ $:=\frac{1}{r^{2}}\|u(r,\cdot)\|_{H^{2}}^{2}+\|u(r,\cdot)\|_{H^{1}}^{2}+\|v(r,\cdot)\|_{H^{1}}^{2}.$ (4)
In the $r$-dependent function space $X_{r}$, the dynamics of the asymptotic hyperbolic operator dominates norm estimates as $r$ goes to infinity while the influence of the spherical Laplacian remains bounded under the weighted $H^{1}$ norm. This choice is the key trick to applying robustness in this setting and allows us to prove our result.
Overview of Paper. In section 2, we present this framework for the existence of a far-field dichotomy for (2) in greater detail. Later, in Section 3, we prove the Lemmas and existence results stated in Section 2. Finally, in Section 4, we discuss possible applications of our results, particularly with respect to the existence and construction of localized spatial patterns in $\mathbb{R}^{n}$.
2 Definitions and Results.
Our approach to proving the existence of exponential dichotomies relies heavily on the property of robustness, the persistence of exponential dichotomies through sufficiently small perturbations. Therefore, to prove that (2) has an exponential dichotomy on an appropriately chosen function space, we decompose it into a principal part that we can directly prove has an exponential dichotomy and a perturbation that we can bound appropriately.
Often the asymptotic operator, $\mathcal{A}_{\infty}=\lim_{r\to\infty}\mathcal{A}(r)$, is a suitable choice for the principal part of a robustness argument. Finding a principal operator is not so simple in this case, however. Assuming that $V(r,\varphi)$ decays to a constant potential $V_{\infty}\in\mathbb{C}^{d\times d}$ as $r$ goes to infinity, the asymptotic operator corresponding to (2) is given by
$\mathcal{A}_{\infty}=\begin{pmatrix}0&1\\ V_{\infty}&0\end{pmatrix}$ (5)
and is associated with the asymptotic equation:
$\mathbf{u}_{r}=\mathcal{A}_{\infty}\mathbf{u}.$ (6)
If we were to take the above $\mathcal{A}_{\infty}$ as the principal operator, we would be left with a remainder that is unbounded relative to $\mathcal{A}_{\infty}$ under the natural product norm on $X$ due to the spherical Laplacian term. Since the spherical Laplacian term can therefore not be left as a perturbation, we must include it in our principal operator and instead choose a norm that bounds the Laplacian when $r$ is small and highlights the effects of $V_{\infty}$ when $r$ is large. This is accomplished by the norms defined in (4). Notice that this choice of an $r$-dependent norm effectively creates a family of function spaces $X_{r}$ where the subscript denotes which value of $r$ is to be used in the norm. [^1] Recall that $X=H^{1}(S^{n-1},\mathbb{C}^{d})\times L^{2}(S^{n-1},C^{d})$. Consequently, a function $u:S^{n-1}\to\mathbb{C}^{d}$ may live in any $X_{r}$, however we will typically consider $u:[0,\infty)\times S^{n-1}\to\mathbb{C}^{d}$ evaluated at a fixed value of $r$ as living in the corresponding $X_{r}$ space; for example, $u(r,\cdot)\in X_{r}$ and $u(\rho,\cdot)\in X_{\rho}$.
With this new weighted family of function spaces, $X_{r}$, we can prove the existence of an exponential dichotomy for a principal part containing the leading order derivatives. This principal part is given by
$\tilde{\mathcal{A}}(r)=\begin{pmatrix}0&1\\ V_{\infty}-\frac{1}{r^{2}}\Delta_{S^{n-1}}&0\end{pmatrix}$ (7)
and leaves a remaining perturbation given by
$\tilde{B}(r)=\begin{pmatrix}0&0\\ V(r,\varphi)-V_{\infty}&-\frac{(n-1)}{r}\end{pmatrix}$. (8)
We will later refer to the principal equation:
$\mathbf{u}_{r}=\tilde{\mathcal{A}}(r)\mathbf{u}.$ (9)
Before further discussing the existence proof, we must first define what is meant by an exponential dichotomy on a family of functions spaces such as $X_{r}$. To provide a solid foundation for this definition, we first state the standard definition of an exponential dichotomy on a general, fixed, Banach space $Y$.
Definition 2.1.
We say that a radial spatial evolutionary equation of the form $\mathbf{u}_{r}=\mathcal{L}(r)\mathbf{u}$ has an exponential dichotomy in a Banach space $Y$ on a sub-interval $J\subseteq\mathbb{R}^{+}$ if there exist positive constants $C$ and $\eta$, two strongly continuous family of projections $P^{\mathrm{s}/\mathrm{u}}:J\rightarrow L(Y)$ with $P^{\mathrm{s}}(r)+P^{\mathrm{u}}(r)=I$, and two families of linear operators $\Phi^{\mathrm{s}/\mathrm{u}}(r,\rho)$ such that the following statements are true:
  • (Stability) For all $\rho\in J$ and $\mathbf{u}_{0}\in Y$ there exists a solution $\Phi^{\mathrm{s}}(r,\rho)\mathbf{u}_{0}$, defined and continuous for both $r$ and $\rho$ on $r\geq\rho$, and differentiable in both $r$ and $\rho$ for $r>\rho$. The operator $P^{\mathrm{s}}(\rho)\mathbf{u}_{0}:=\Phi^{\mathrm{s}}(\rho,\rho)\mathbf{u}_{0}$ gives the stable projection, and
    $\|\Phi^{\mathrm{s}}(r,\rho)\mathbf{u}_{0}\|_{Y}\leq Ce^{-\eta(r-\rho)}\|\mathbf{u}_{0}\|_{Y},\qquad r\geq\rho,\,\,r,\rho\in J.$ (10)
  • (Instability) For all $\rho\in J$ and $\mathbf{u}_{0}\in Y$ there exists a solution $\Phi^{\mathrm{u}}(r,\rho)\mathbf{u}_{0}$, defined and continuous for both $r$ and $\rho$ on $r\leq\rho$, and differentiable in both $r$ and $\rho$ for $r<\rho$. The operator $P^{\mathrm{u}}(\rho)\mathbf{u}_{0}:=\Phi^{\mathrm{u}}(\rho,\rho)\mathbf{u}_{0}$ gives the unstable projection, and
    $\|\Phi^{\mathrm{u}}(r,\rho)\mathbf{u}_{0}\|_{Y}\leq Ce^{\eta(r-\rho)}\|\mathbf{u}_{0}\|_{Y},\qquad r\leq\rho,\,\,r,\rho\in J.$ (11)
  • (Invariance) The solutions $\Phi^{\mathrm{s}}(r,\rho)\mathbf{u}_{0}$ and $\Phi^{\mathrm{u}}(r,\rho)\mathbf{u}_{0}$ are invariant under the respective projections. That is, $\Phi^{\mathrm{s}}(r,\rho)\mathbf{u}_{0}\in\mathrm{Rg}\,P^{\mathrm{s}}(r)$ for all $r\geq\rho\in J$ and $\Phi^{\mathrm{u}}(r,\rho)\mathbf{u}_{0}\in\mathrm{Rg}\,P^{\mathrm{u}}(r)$ for all $r\leq\rho\in J$. Equivalently,
    $~P^{\mathrm{s}}(r)\Phi^{\mathrm{s}}(r,\rho)=\Phi^{\mathrm{s}}(r,\rho)=\Phi^{\mathrm{s}}(r,\rho)P^{\mathrm{s}}(\rho)$
    $P^{\mathrm{u}}(r)\Phi^{\mathrm{u}}(r,\rho)=\Phi^{\mathrm{u}}(r,\rho)=\Phi^{\mathrm{u}}(r,\rho)P^{\mathrm{u}}(\rho).$
Since $X_{r}$ is a family of Banach spaces whose norm changes with $r$, we now define an analogous understanding of an exponential dichotomy on $X_{r}$.
Definition 2.2.
(Definition 5.2 in [17]) We say that a radial spatial evolutionary equation of the form $\mathbf{u}_{r}=\mathcal{L}(r)\mathbf{u}$ has an exponential dichotomy in the $r$-dependent family of Banach spaces $X_{r}$ on a sub-interval $J\subseteq\mathbb{R}^{+}$ if there exist positive constants $C$ and $\eta$, two strongly continuous family of projections $P^{\mathrm{s}/\mathrm{u}}(r)\in L(X_{r})$ with $P^{\mathrm{s}}(r)+P^{\mathrm{u}}(r)=I$, and two families of linear operators $\Phi^{\mathrm{s}/\mathrm{u}}(r,\rho)\in L(X_{\rho},X_{r})$ such that the following statements are true:
  • (Stability) For all $\rho\in J$ and $\mathbf{u}_{0}\in X_{\rho}$ there exists a solution $\Phi^{\mathrm{s}}(r,\rho)\mathbf{u}_{0}$, defined and continuous for both $r$ and $\rho$ on $r\geq\rho$, and differentiable in both $r$ and $\rho$ for $r>\rho$. The operator $P^{\mathrm{s}}(\rho)\mathbf{u}_{0}:=\Phi^{\mathrm{s}}(\rho,\rho)\mathbf{u}_{0}$ gives the stable projection, and
    $\|\Phi^{\mathrm{s}}(r,\rho)\mathbf{u}_{0}\|_{X_{r}}\leq Ce^{-\eta(r-\rho)}\|\mathbf{u}_{0}\|_{X_{\rho}},\qquad r\geq\rho,\,\,r,\rho\in J.$ (12)
  • (Instability) For all $\rho\in J$ and $\mathbf{u}_{0}\in X_{\rho}$ there exists a solution $\Phi^{\mathrm{u}}(r,\rho)\mathbf{u}_{0}$, defined and continuous for both $r$ and $\rho$ on $r\leq\rho$, and differentiable in both $r$ and $\rho$ for $r<\rho$. The operator $P^{\mathrm{u}}(\rho)\mathbf{u}_{0}:=\Phi^{\mathrm{u}}(\rho,\rho)\mathbf{u}_{0}$ gives the unstable projection, and
    $\|\Phi^{\mathrm{u}}(r,\rho)\mathbf{u}_{0}\|_{X_{r}}\leq Ce^{\eta(r-\rho)}\|\mathbf{u}_{0}\|_{X_{\rho}},\qquad r\leq\rho,\,\,r,\rho\in J.$ (13)
  • (Invariance) The solutions $\Phi^{\mathrm{s}}(r,\rho)\mathbf{u}_{0}$ and $\Phi^{\mathrm{u}}(r,\rho)\mathbf{u}_{0}$ are invariant under the respective projections. That is, $\Phi^{\mathrm{s}}(r,\rho)\mathbf{u}_{0}\in\mathrm{Rg}\,P^{\mathrm{s}}(r)$ for all $r\geq\rho\in J$ and $\Phi^{\mathrm{u}}(r,\rho)\mathbf{u}_{0}\in\mathrm{Rg}\,P^{\mathrm{u}}(r)$ for all $r\leq\rho\in J$. Equivalently,
    $~P^{\mathrm{s}}(r)\Phi^{\mathrm{s}}(r,\rho)=\Phi^{\mathrm{s}}(r,\rho)=\Phi^{\mathrm{s}}(r,\rho)P^{\mathrm{s}}(\rho)$
    $P^{\mathrm{u}}(r)\Phi^{\mathrm{u}}(r,\rho)=\Phi^{\mathrm{u}}(r,\rho)=\Phi^{\mathrm{u}}(r,\rho)P^{\mathrm{u}}(\rho).$
For completeness, we now also state the standard robustness result for exponential dichotomies. This result was originally stated for an arbitrary fixed infinite dimensional function space $Y$ in [11], however, the statement holds almost identically for the family of spaces $X_{r}$ [17]. Here we summarize both robustness statements in one.
Proposition 2.3.
(Theorem 1 in [11] and Proposition 5.9 in [17]) Exponential dichotomies are robust under sufficiently small perturbations. More precisely, if a system of the form $\mathbf{u}_{r}=\mathcal{L}(r)\mathbf{u}$ has an exponential dichotomy on some interval $J\subseteq\mathbb{R}^{+}$ in $Y$ (or in the $r$-dependent family of $X_{r}$ spaces) with constants $C,\eta>0$, then for all $\varepsilon>0$ and $\tilde{\eta}\in(0,\eta)$ there are constants $\tilde{C},\delta>0$ such that the perturbed system
$\mathbf{u}_{r}=(\mathcal{L}(r)+\mathcal{B}(r))\mathbf{u}$
with
$\|\mathcal{B}(r)\|_{L(Y)}\leq\delta,\quad\qquad\big(\text{ or }\quad\|\mathcal{B}(r)\|_{L(X_{r})}\leq\delta,\big)\qquad r\in J$
has an exponential dichotomy on $J$ in the same (family of) function space(s) with constants $\tilde{C}$ and $\tilde{\eta}$ and with projections $\varepsilon$-close to the projections of $\mathbf{u}_{r}=\mathcal{L}(r)\mathbf{u}$. Furthermore, if $\|\mathcal{B}(r)\|_{L(Y)}=O(1/r)$ (or $\|\mathcal{B}(r)\|_{L(X_{r})}=O(1/r)$), then one can take $\tilde{\eta}=\eta$.
With these robustness results in hand, our aim is to prove the existence of an exponential dichotomy for the principal system (9) given the existence of an exponential dichotomy for the asymptotic system (6). Assuming that $V(r,\varphi)$ decays to $V_{\infty}$ sufficiently fast, Proposition 2.3 can then be applied directly to establish our main result. We therefore make the following hypothesis:
(H1.) $V(r,\varphi)$ decays to $V_{\infty}$ sufficiently fast as $r\to\infty$. That is, $\|V(r,\varphi)-V_{\infty}\|\leq C/r$ for large $r$.
Since $\mathcal{A}_{\infty}$ is autonomous and homogeneous across $S^{n-1}$, the asymptotic system (6) has an exponential dichotomy whenever $\mathcal{A}_{\infty}$ is hyperbolic as a $2d\times 2d$ matrix. We can construct an exponential dichotomy for (6) via a finite spectral decomposition whenever the eigenvalues of $\mathcal{A}_{\infty}$ are bounded away from $i\mathbb{R}\subset\mathbb{C}$ or, equivalently, whenever the eigenvalues of $V_{\infty}$ are bounded away from the negative real line, $(-\infty,0]$. To use this condition later, we now set it aside in a hypothesis:
(H2.) All eigenvalues of $V_{\infty}$ are bounded away from the negative real line $(-\infty,0]$.
Finally, we also assume that $V_{\infty}$ has only semi-simple eigenvalues, and is therefore diagonalizable. This last hypothesis greatly simplifies the proofs to follow, allowing for an explicit spectral decomposition of $\tilde{\mathcal{A}}(r)$ into scalar pieces as we will see in §3.1. However, similar arguments should hold in the case of nontrivial Jordan blocks.
(H3.) $V_{\infty}$ is diagonalizable with eigenvalues $\{\lambda_{\ell}\}_{\ell=1}^{d}$ and corresponding eigenvectors $\{\rho_{\ell}\}_{\ell=1}^{d}\subset\mathbb{C}^{d}$.
Hypotheses (H2) and (H3) allow us to fully characterize an exponential dichotomy for the asymptotic system (6). We must then understand the relationship between the principal operator and the asymptotic operator. To do this, we establish an isomorphism, or rather a family of isomorphisms, that can translate between the $X_{r}$ setting and the $L^{2}\!\!\times\!L^{2}$ setting. As we will see in Lemma 2.4, the basis of this family of isomorphisms is the differential operator defined by
$A(r):=V_{\infty}-\frac{1}{r^{2}}\Delta_{S^{n-1}}.$ (14)
Note that with this new notation, the principal operator (7) can be written as $\tilde{\mathcal{A}}(r)=\left(\begin{smallmatrix}0&1\\ A(r)&0\end{smallmatrix}\right)$.
Lemma 2.4.
If (H2) and (H3) hold then the following statements are true:
  1. (i)
    The family of operators $A(r)^{1/2}:H^{1}(S^{n-1})\rightarrow L^{2}(S^{n-1})$ is well-defined, bounded uniformly in $r$, strongly continuous, and continuously differentiable in $r$, with
    $\left\|\partial_{r}\!\left[A(r)^{1/2}\right]\!A(r)^{-1/2}\right\|_{L(L^{2})}\leq Cr^{-1}$
    for some constant $C>0$ independent of $r$.
  2. (ii)
    For each $r\geq 1$, the operator $(u,v)\mapsto(A(r)^{1/2}u,v)$ defined on $X_{r}\rightarrow L^{2}(S^{n-1})\times L^{2}(S^{n-1})$ is an isomorphism and satisfies
    $C_{1}\left(\|A(r)^{1/2}u\|_{L^{2}}^{2}+\|v\|_{L^{2}}^{2}\right)\leq\|\mathbf{u}\|_{X_{r}}^{2}\leq C_{2}\left(\|A(r)^{1/2}u\|^{2}_{L^{2}}+\|v\|_{L^{2}}^{2}\right)$ (15)
    for all $\mathbf{u}=(u,v)\in X_{r}$ with constants $C_{1},C_{2}>0$ independent of both $r$ and $\mathbf{u}$.
Lemma 2.4 establishes $(u,v)\mapsto(A(r)^{1/2}u,v)$ as an isomorphism from $X_{r}$ to $L^{2}\!\!\times\!L^{2}$. Applying the corresponding change of variables $\tilde{u}=A(r)^{1/2}u$ to the principal equation (9) we get the following equation posed on $L^{2}\!\!\times\!L^{2}$:
$\mathbf{\tilde{u}}_{r}=\begin{pmatrix}\tilde{u}_{r}\\ v_{r}\end{pmatrix}=\begin{pmatrix}A(r)^{1/2}v+\partial_{r}\!\left[A(r)^{1/2}\right]A(r)^{-1/2}\tilde{u}\\ A(r)^{1/2}\tilde{u}\end{pmatrix}=\begin{pmatrix}0&A(r)^{1/2}\\ A(r)^{1/2}&0\end{pmatrix}\mathbf{\tilde{u}}+\mathcal{O}\left(1/r\right)\mathbf{\tilde{u}}$ (16)
By Proposition 2.3 and Lemma 2.4, an exponential dichotomy for (2) in $X_{r}$ can be established by proving the existence of an exponential dichotomy for the principal part of (16) in $L^{2}\!\!\times\!L^{2}$.
We proceed by performing another change of variables, $w^{\pm}=\tilde{u}\pm v$, to diagonalize the remaining equation:
$\mathbf{w}_{r}=\begin{pmatrix}w^{+}_{r}\\ w^{-}_{r}\end{pmatrix}=\begin{pmatrix}A(r)^{1/2}w^{+}\\ -A(r)^{1/2}w^{-}\end{pmatrix}=\begin{pmatrix}A(r)^{1/2}&0\\ 0&-A(r)^{1/2}\end{pmatrix}\mathbf{w}.$ (17)
Once we arrive at (17), it is relatively straightforward to construct an exponential dichotomy using the spectral decomposition of the spherical Laplacian and the finite spectral decomposition of $V_{\infty}$ (diagonal thanks to our hypothesis (H3)).
Lemma 2.5.
Suppose that (H1), (H2), and (H3) all hold. Then equation (17) has an exponential dichotomy,  $\tilde{\Phi}^{\mathrm{s}/\mathrm{u}}(r,t)$, on $L^{2}(S^{n-1})\times L^{2}(S^{n-1})$ with the same decay rate as the asymptotic equation (6):  $\eta=\displaystyle{\tilde{\eta}=\min_{1\leq\ell\leq d}\mathrm{Re}[\lambda_{\ell}^{1/2}]}$.
This provides the last stepping stone before our primary result:
Theorem 2.1.
If (H1), (H2), and (H3) all hold, then (2) has an exponential dichotomy in $X_{r}$ for $r>1$ with the same decay rate as the asymptotic equation (6): ${\eta=\tilde{\eta}=\min_{1\leq\ell\leq d}\mathrm{Re}[\lambda_{\ell}^{1/2}]}$.
3 Proofs.
In this section we will fill in the details of the argument by proving Lemma 2.4, Lemma 2.5, and Theorem 2.1. We begin by discussing, in detail, the spectral decomposition of the spherical Laplacian and the resulting multiplication symbols for the operators $A(r)$ and $A(r)^{1/2}$.
3.1 Spectral Decomposition via Spherical Harmonics.
For each degree $k\geq 0$ with $k\in\mathbb{Z}$, there is an orthonormal set $\{Y_{k}^{j}\}_{j=1}^{\nu(k)}\subset L^{2}(S^{n-1},\mathbb{C})$ which spans a $\nu(k)$-dimensional eigenspace and satisfies
$\Delta_{S^{n-1}}Y_{k}^{j}=\mu_{k}Y_{k}^{j}\quad\text{where}\quad\mu_{k}=k(k+n-2).$
These $\{Y_{k}^{j}\}$ are the spherical harmonics of $\Delta_{S^{n-1}}$, and for each $(u(r,\cdot),v(r,\cdot))\in L^{2}(S^{n-1},\mathbb{C}^{d})\times L^{2}(S^{n-1},\mathbb{C}^{d})$ they allow a spectral decomposition given by
$u(r,\varphi)$ $=\sum_{k=0}^{\infty}\sum_{j=1}^{\nu(k)}u_{kj}(r)Y_{k}^{j}(\varphi)\quad\text{where}\quad u_{kj}(r)=\int_{S^{n-1}}u(r,\varphi)\overline{Y_{k}^{j}(\varphi)}d\varphi\in\mathbb{C}^{d}$
$\text{and}\quad v(r,\varphi)$ $=\sum_{k=0}^{\infty}\sum_{j=1}^{\nu(k)}v_{kj}(r)Y_{k}^{j}(\varphi)\quad\text{where}\quad v_{kj}(r)=\int_{S^{n-1}}v(r,\varphi)\overline{Y_{k}^{j}(\varphi)}d\varphi\in\mathbb{C}^{d}.$
This decomposition also allows for the definition of alternative, equivalent Sobolev norms on $H^{\alpha}(S^{n-1},\mathbb{C}^{d})$, and similarly on $L^{2}(S^{n-1},\mathbb{C}^{d})$:
$\|u(r,\cdot)\|_{H^{\alpha}}^{2}:=\sum_{k=0}^{\infty}\sum_{j=1}^{\nu(k)}(1+\mu_{k})^{\alpha}\|u_{kj}(r)\|^{2}_{\mathbb{C}^{d}},\qquad\|u(r,\cdot)\|_{L^{2}}^{2}:=\sum_{k=0}^{\infty}\sum_{j=1}^{\nu(k)}\|u_{kj}(r)\|_{\mathbb{C}^{d}}^{2}$ (18)
According to this decomposition, Equation (9) is reduced in each degree $k$ (suppressing the $j$ subscript as each equation is the same) to
$\begin{pmatrix}{u_{k}}^{\prime}\\ {v_{k}}^{\prime}\end{pmatrix}=\begin{pmatrix}0&1\\ V_{\infty}+\mu_{k}/r^{2}&0\end{pmatrix}\begin{pmatrix}u_{k}\\ v_{k}\end{pmatrix}=\begin{pmatrix}0&1\\ A_{k}(r)&0\end{pmatrix}\begin{pmatrix}u_{k}\\ v_{k}\end{pmatrix}$ (19)
where $A_{k}(r)=V_{\infty}+\mu_{k}/r^{2}\in\mathbb{C}^{d\times d}$ and $(u_{k}(r),v_{k}(r))\in\mathbb{C}^{d}\times\mathbb{C}^{d}$.
Notice that this spectral decomposition reveals $A(r)$ as a multiplication operator with symbol $A_{k}(r)$. Similarly, $A(r)^{1/2}$ can be considered a multiplication operator with symbol $A_{k}(r)^{1/2}$. When $d>1$, we can further decompose the matrices $A_{k}(r)$ and $A_{k}(r)^{1/2}$ in terms of the Jordan normal form of $V_{\infty}$. Recalling that by hypothesis (H3), $V_{\infty}$ is diagonalizable with eigenvectors $\{\rho_{1},\cdots,\rho_{d}\}$, let $R$ be the matrix whose columns are each $\rho_{\ell}$. Then both $A_{k}(r)$ and $A_{k}(r)^{1/2}$ are also diagonalizable with the same eigenvectors as $V_{\infty}$:
$A_{k}(r)$ $=R\begin{pmatrix}\lambda_{1}&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\!\lambda_{d}\end{pmatrix}R^{-1}+\mu_{k}/r^{2}=R\begin{pmatrix}\lambda_{1}+\mu_{k}/r^{2}&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\lambda_{d}+\mu_{k}/r^{2}\end{pmatrix}R^{-1}\quad\text{and}$
$A_{k}(r)^{1/2}$ $=R\begin{pmatrix}\sqrt{\lambda_{1}+\mu_{k}/r^{2}}&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\sqrt{\lambda_{d}+\mu_{k}/r^{2}}\end{pmatrix}R^{-1}=R\begin{pmatrix}\gamma_{k1}(r)&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\gamma_{kd}(r)\end{pmatrix}R^{-1}$
where $\gamma_{k\ell}(r):=\sqrt{\lambda_{\ell}+\mu_{k}/r^{2}}$ are the eigenvalues of $A_{k}(r)^{1/2}$. In the case of nontrivial Jordan blocks, a similar decomposition of the symbols $A_{k}(r)$ and $A_{k}(r)^{1/2}$ would still reveal each $\gamma_{k\ell}(r)$ to be an eigenvalue of $A_{k}(r)^{1/2}$ with multiplicity inherited from $V_{\infty}$. In fact, the invariant generalized eigenspace of each eigenvalue would be preserved though the exact generalized eigenvectors would not.
In the analysis that follows, we will use the eigenvector basis $R=[\rho_{1},\ldots,\rho_{d}]$ for $\mathbb{C}^{d}$ to decompose $A(r)$, $A(r)^{1/2}$, and other related operators into scalar pieces. In particular, we will use this decomposition and the equivalence of finite norms to bound the $L^{2}$ norm of $A_{k}(r)^{1/2}$ and related operators in the proof of Lemma 2.4. This looks like the following:
$\left\|A(r)^{1/2}u\right\|_{L^{2}}^{2}=\sum_{k,j}$ $\left\|A_{k}(r)^{1/2}u_{kj}\right\|^{2}_{\mathbb{C}^{d}}=\sum_{k,j}\left\|R\begin{pmatrix}\gamma_{k1}(r)&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\gamma_{kd}(r)\end{pmatrix}R^{-1}u_{kj}\right\|^{2}_{\mathbb{C}^{d}}$
$\leq\sum_{k,j}\left\|R\right\|^{2}\left(\max_{1\leq\ell\leq d}|\gamma_{k\ell}(r)|^{2}\right)\left\|R^{-1}\right\|^{2}\left\|u_{kj}\right\|_{\mathbb{C}^{d}}^{2}\leq C\max_{1\leq\ell\leq d}|\gamma_{k\ell}(r)|^{2}\|u\|_{L^{2}}^{2}$ (20)
We are now ready to proceed with the proof of Lemma 2.4 keeping in mind both the bound given by Equation (20) and the technique used to derive it. Note that a similar decomposition, not into scalar pieces but into nilpotent blocks, should yield a parallel argument if we relax hypothesis (H3) to consider nontrivial Jordan blocks in the spectral structure of $V_{\infty}$.
3.2 Proofs of Lemmas 2.4 and 2.5.
In the proofs that follow we rely on a series of small lemmas which are stated and proven in the Appendix. We also reference Equation (20) to emphasize the technique used to rewrite an $L^{2}$ norm into scalar factors.
Proof (of Lemma 2.4)..
We prove statement (i) in five steps:
  1. 1.
    $A(r)^{1/2}$ is well-defined and uniformly bounded in $r$.
    Notice that there always exists a constant $K>0$, independent of both $\lambda_{\ell}$ and $r>1$, such that $|\gamma_{k\ell}(r)^{2}|=|\lambda_{\ell}+\mu_{k}/r^{2}|\leq K(1+\mu_{k})$. Then, by Equation (20), we have that
    $\|A(r)^{1/2}u\|_{L^{2}}^{2}\leq\sum_{k,j}C\max_{1\leq\ell\leq d}|\gamma_{k\ell}(r)|^{2}\|u_{kj}\|^{2}_{\mathbb{C}^{d}}\leq\sum_{k,j}\tilde{C}(1+\mu_{k})\|u_{kj}\|^{2}_{\mathbb{C}^{d}}=\tilde{C}\|u\|_{H^{1}}^{1}.$
  2. 2.
    $A(r)^{1/2}$ is strongly continuous.
    Let $r_{1},r_{2}\geq 1$ be sufficiently close so that $r_{1}\leq 2r_{2}$. Then
    $\|A(r_{1})^{1/2}u-A(r_{2})^{1/2}u\|_{L^{2}}^{2}$ $\leq\sum_{k,j}C\max_{1\leq\ell\leq d}|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})|^{2}\|u_{k,j}\|^{2}_{\mathbb{C}^{d}}$ by Eq (20)
[…]
  3. 3.
    $A(r)^{1/2}$ is differentiable.
    We claim that the operator $\partial_{r}\!\left[A(r)^{1/2}\right]$, defined via the symbol $\partial_{r}\gamma_{k\ell}(r)=\left(\frac{-\mu_{k}/r^{3}}{\gamma_{k\ell}(r)}\right)$, is the Fréchet derivative of $A(r)^{1/2}$. To show this, fix $r\geq 1$ and let $r^{\prime}\geq 1$ be sufficiently close to $r$ so that $r^{\prime}/2\leq r<2r^{\prime}$. Then
    $\|A(r)^{1/2}u-A(r^{\prime})^{1/2}u$ $-\partial_{r}A(r)^{1/2}(r-r^{\prime})u\|_{L^{2}}^{2}$
    $\leq\sum_{k,j}C\max_{1\leq\ell\leq d}|\gamma_{k\ell}(r)-\gamma_{k\ell}(r^{\prime})-\partial_{r}\gamma_{k\ell}(r)(r-r^{\prime})|^{2}\|u_{k,j}\|^{2}_{\mathbb{C}^{d}}$ by Eq (20)
[…]
    This leads us to conclude that $\partial_{r}A(r)^{1/2}$ satisfies the limit definition of the Fréchet derivative:
    $\lim_{r^{\prime}\to r}\frac{\|A(r)^{1/2}-A(r^{\prime})^{1/2}-\partial_{r}A(r)^{1/2}(r-r^{\prime})\|_{L^{2}}}{|r-r^{\prime}|}\leq\lim_{r^{\prime}\to r}\frac{C^{\prime}|r-r^{\prime}|^{2}}{|r-r^{\prime}|}=0.$
  4. 4.
    $\partial_{r}A(r)^{1/2}$ is continuous.
    Let $r_{1}\geq r_{2}\geq 1$ be sufficiently close so that $r_{2}/2\leq r_{1}\leq 2r_{2}$. Then
    $\|\partial_{r}A(r_{1})^{1/2}u-\partial_{r}A(r_{2})^{1/2}u\|_{L^{2}}^{2}$ $\leq\sum_{k,j}C\max_{1\leq\ell\leq d}|\partial_{r}\gamma_{k\ell}(r_{1})-\partial_{r}\gamma_{k\ell}(r_{2})|^{2}\|u_{kj}\|^{2}_{\mathbb{C}^{d}}$ by Eq (20)
    $\leq\tilde{C}\left||r_{1}-r_{2}|+|r_{1}^{3}-r_{2}^{3}|\right|^{2}\sum_{k,j}(1+\mu_{k})\|u_{k,j}\|^{2}_{\mathbb{C}^{d}}\quad$ by Lemma A.6
    $\leq\tilde{C}\left(|r_{1}-r_{2}|+|r_{1}^{3}-r_{2}^{3}|\right)^{2}\|u\|_{H^{1}}^{2}.$
  5. 5.
    The operator norm of $\partial_{r}\!\left[A(r)^{1/2}\right]\!A(r)^{-1/2}:L^{2}\to L^{2}$ is of order $1/r$.
    First note that the symbol of $A(r)^{-1/2}$ is given by $\gamma_{k\ell}(r)^{-1}$ and that consequently the symbol of the product $\partial_{r}\!\left[A(r)^{1/2}\right]\!A(r)^{-1/2}$ is given by $\left(\frac{-\mu_{k}/r^{3}}{\gamma_{k\ell}(r)^{2}}\right)$. We therefore obtain a bound on the $L^{2}$ operator norm of $\partial_{r}\!\left[A(r)^{1/2}\right]\!A(r)^{-1/2}$ by bounding the supremum of the symbols over all values of $k$ and $\ell$. This bound, given by
    $\sup_{k,\ell}\left|\frac{-\mu_{k}/r^{3}}{\gamma_{k\ell}(r)^{2}}\right|=\left(\frac{1}{r}\right)\sup_{k,\ell}\frac{|\mu_{k}/r^{2}|}{|\lambda_{\ell}+\mu_{k}/r^{2}|}\leq Cr^{-1},$
    exists with $C$ independent of $r$ by hypothesis (H2.). Consequently, by Equation (20),
    $\left\|\partial_{r}\!\left[A(r)^{1/2}\right]\!A(r)^{-1/2}u\right\|_{L^{2}}^{2}$ $\leq\sum_{k,j}C\max_{1\leq\ell\leq d}\left|\frac{-\mu_{k}/r^{3}}{\gamma_{k\ell}(r)^{2}}\right|\|u_{kj}(r)\|_{\mathbb{C}^{d}}^{2}\leq\tilde{C}r^{-1}\sum_{kj}\|u_{kj}\|_{\mathbb{C}^{d}}^{2}=\tilde{C}r^{-1}\|u\|_{L^{2}}^{2}$
    with $\tilde{C}$ also independent of $r$, and the operator norm on $\partial_{r}\!\left[A(r)^{1/2}\right]\!A(r)^{-1/2}$ is $O(r^{-1})$.
To prove (ii), we first prove that $\|A(r)^{1/2}u\|_{L^{2}}^{2}+\|v\|_{L^{2}}^{2}$ is bounded above by a constant times $\|\mathbf{u}\|_{X_{r}}^{2}$:
$\|A(r)^{1/2}u\|_{L^{2}}^{2}+\|v\|_{L^{2}}^{2}$ $\leq\sum_{k,j}C\max_{1\leq\ell\leq d}\left\{\left(|\lambda_{\ell}|+\mu_{k}/r^{2}\right)\right\}\|u_{k,j}\|^{2}_{\mathbb{C}^{d}}+\|v\|_{L^{2}}^{2}$ by Eq (20)
[…]
On the other hand,
$\|\mathbf{u}\|_{X_{r}}^{2}$ $=\frac{1}{r^{2}}\|u\|_{H^{1}}^{2}+\|u\|_{L^{2}}^{2}+\|v\|_{L^{2}}^{2}$
[…]
This concludes the proof that $A(r)^{1/2}$ yields an isomorphism between $X_{r}$ and $L^{2}\times L^{2}$. ∎
The proof of Lemma 2.5 uses the same spectral decomposition ideas to construct an exponential dichotomy for (17) from the hyperbolicity of $A_{\infty}$ given by (H2).
Proof (of Lemma 2.5)..
Let $\tilde{\Phi}^{\mathrm{s}/\mathrm{u}}$ be defined as multiplication operators as follows:
$\tilde{\Phi}^{\mathrm{s}}(r,t)\mathbf{w}$ $=\begin{pmatrix}0&0\\ 0&\mathrm{exp}\left(-\int_{t}^{r}A(s)^{1/2}ds\right)\end{pmatrix}\mathbf{w}=\sum_{k,j}\begin{pmatrix}0&0\\ 0&\mathrm{exp}\left(-\int_{t}^{r}A_{k}(s)^{1/2}ds\right)\end{pmatrix}\mathbf{w}_{kj}Y_{k}^{j}(\varphi),\quad r\geq t$
[…]
where by the spectral structure of $V_{\infty}$, we can further decompose (21) according to
$\mathrm{exp}\left(\pm\int_{t}^{r}A_{k}(s)^{1/2}ds\right)=R\begin{pmatrix}\mathrm{exp}\left(\pm\int_{t}^{r}\gamma_{k1}(s)^{1/2}ds\right)&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\mathrm{exp}\left(\pm\int_{t}^{r}\gamma_{kd}(s)^{1/2}ds\right)\end{pmatrix}R^{-1}.$ (22)
The continuity and differentiability of (21) follows from the continuity and differentiability of both $A(r)^{1/2}$ and each $\gamma_{k\ell}(r)$ (Lemmas 2.4, A.4, A.5, and A.6). By taking the derivatives of $\tilde{\Phi}^{\mathrm{s}}(r,t)\mathbf{w}$ and $\tilde{\Phi}^{\mathrm{u}}(r,t)\mathbf{w}$ with respect to $r$ one can directly confirm that for each $t\in\mathbb{R}$ and $\mathbf{w}\in(L^{2}(S^{n-1})\times L^{2}(S^{n-1}))$ these are both solutions to (17). The invariance conditions of Definition 2.1 are also clearly satisfied by the evolutionary operators and their corresponding projections, $\tilde{P}^{\mathrm{s}}:=\tilde{\Phi}^{\mathrm{s}}(r,r)=\left(\begin{smallmatrix}0&0\\ 0&1\end{smallmatrix}\right)$ and $\tilde{P}^{\mathrm{u}}:=\tilde{\Phi}^{\mathrm{u}}(r,r)=\left(\begin{smallmatrix}1&0\\ 0&0\end{smallmatrix}\right)$.
It remains to be shown that the exponential decay estimates hold for $\tilde{\eta}:=\min_{1\leq\ell\leq d}\mathrm{Re}[\lambda_{\ell}^{1/2}]$. These decay estimates follow from the fact that for all $k\in\{0,1,2,\ldots\}$ and $r>0$, $\mathrm{Re}\left[\gamma_{k\ell}(r)\right]\geq\mathrm{Re}[\lambda_{\ell}^{1/2}]$ (see Lemma A.8). We therefore obtain:
$\left\|\tilde{\Phi}^{s}(r,t)\mathbf{w}\right\|_{L^{2}\times L^{2}}^{2}$ $=\sum_{k,j}\left\|\mathrm{exp}\left(-\int_{t}^{r}A_{k}(s)^{1/2}ds\right)w^{-}_{kj}\right\|_{\mathbb{C}^{d}}^{2}$
$=\sum_{k,j}\left\|R\begin{pmatrix}\mathrm{exp}\left(-\int_{t}^{r}\gamma_{k1}(s)^{1/2}ds\right)&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\mathrm{exp}\left(-\int_{t}^{r}\gamma_{kd}(s)^{1/2}ds\right)\end{pmatrix}R^{-1}w^{-}_{kj}\right\|_{\mathbb{C}^{d}}^{2}$
$\leq\sum_{k,j}\|R\|^{2}\left(\max_{1\leq\ell\leq d}\left|\mathrm{exp}\left(-\int_{t}^{r}\gamma_{k\ell}(s)ds\right)\right|^{2}\right)\|R^{-1}\|^{2}\|w^{-}_{kj}\|^{2}_{\mathbb{C}^{d}}$
$\leq\sum_{k,j}\|R\|^{2}\|R^{-1}\|^{2}\left(\max_{1\leq\ell\leq d}\left[e^{-\mathrm{Re}\big[\lambda_{\ell}^{1/2}\big](r-t)}\right]^{2}\right)\|w^{-}_{kj}\|_{\mathbb{C}^{d}}^{2}$
$\leq\sum_{k,j}C\left(e^{-\min\limits_{1\leq\ell\leq d}2\mathrm{Re}\big[\lambda_{\ell}^{1/2}\big](r-t)}\right)\|w^{-}_{kj}\|_{\mathbb{C}^{d}}^{2}$
$\leq\sum_{k,j}Ce^{-2\tilde{\eta}(r-t)}\|w^{-}_{kj}\|_{\mathbb{C}^{d}}^{2}=Ce^{-2\tilde{\eta}(r-t)}\|w^{-}\|_{L^{2}}^{2}\leq Ce^{-2\tilde{\eta}(r-t)}\|\mathbf{w}\|_{L^{2}\times L^{2}}^{2},$ $r\geq t$
$\left\|\tilde{\Phi}^{u}(r,t)\mathbf{w}\right\|_{L^{2}\times L^{2}}^{2}$ $=\sum_{k,j}\left\|\mathrm{exp}\left(\int_{t}^{r}A_{k}(s)^{1/2}ds\right)w^{+}_{kj}\right\|_{\mathbb{C}^{d}}^{2}$
$=\sum_{k,j}\left\|R\begin{pmatrix}\mathrm{exp}\left(-\int_{r}^{t}\gamma_{k1}(s)^{1/2}ds\right)&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\mathrm{exp}\left(-\int_{r}^{t}\gamma_{kd}(s)^{1/2}ds\right)\end{pmatrix}R^{-1}w^{+}_{kj}\right\|_{\mathbb{C}^{d}}^{2}$
$\leq\sum_{k,j}\|R\|^{2}\left(\max_{1\leq\ell\leq d}\left|\mathrm{exp}\left(-\int_{r}^{t}\gamma_{k\ell}(s)ds\right)\right|^{2}\right)\|R^{-1}\|^{2}\|w^{+}_{kj}\|^{2}_{\mathbb{C}^{d}}$
$\leq\sum_{k,j}\|R\|^{2}\|R^{-1}\|^{2}\left(\max_{1\leq\ell\leq d}\left[e^{-\mathrm{Re}\big[\lambda_{\ell}^{1/2}\big](t-r)}\right]^{2}\right)\|w^{+}_{kj}\|_{\mathbb{C}^{d}}^{2}$
$\leq\sum_{k,j}C\left(e^{-\min\limits_{1\leq\ell\leq d}2\mathrm{Re}\big[\lambda_{\ell}^{1/2}\big](t-r)}\right)\|w^{+}_{kj}\|_{\mathbb{C}^{d}}^{2}$
$\leq\sum_{k,j}Ce^{-2\tilde{\eta}(t-r)}\|w^{+}_{kj}\|_{\mathbb{C}^{d}}^{2}=Ce^{2\tilde{\eta}(r-t)}\|w^{+}\|_{L^{2}}^{2}\leq Ce^{2\tilde{\eta}(r-t)}\|\mathbf{w}\|_{L^{2}\times L^{2}}^{2},$ $r\leq t$
3.3 Proof of Theorem 2.1.
Proof (of Theorem 2.1)..
By Lemma 2.5, (17) has an exponential dichotomy in $L^{2}(S^{n-1})\times L^{2}(S^{n-1})$ for $r>1$ with decay rate $\tilde{\eta}$. Recall that (17) is related to the principal part of (16) via a change of coordinates $w^{\pm}=\tilde{u}\pm v$. This change of coordinates is equivalent to conjugation by the linear transformation $M:=\left(\begin{smallmatrix}1&\,\,\,1\\ 1&-1\end{smallmatrix}\right)$:
$\mathbf{w}_{r}=M\begin{pmatrix}0&A(r)^{1/2}\\ A(r)^{1/2}&0\end{pmatrix}\!M^{-1}\,\mathbf{w}\qquad\text{and}\qquad\mathbf{\tilde{u}}_{r}=M^{-1}\begin{pmatrix}A(r)^{1/2}&0\\ 0&-A(r)^{1/2}\end{pmatrix}\!M\,\mathbf{\tilde{u}}$.
Notice that since $M=2M^{-1}$, the related transformation given by $\tilde{M}:=\tfrac{1}{\sqrt{2}}M$ is a unitary, self-adjoint operator. Consequently, conjugation by $\tilde{M}$ (equivalent to conjugation by $M$) defines an isometric isomorphism on $L^{2}\!\!\times\!L^{2}$, and the principal part of (16) has an exponential dichotomy on $L^{2}\!\!\times\!L^{2}$ for $r>1$ with decay rate $\tilde{\eta}$ defined by
$\hat{\Phi}^{\mathrm{s}}(r,t)$ $:=M^{-1}\tilde{\Phi}^{\mathrm{s}}(r,t)M,\qquad r\geq t$
$\hat{\Phi}^{\mathrm{u}}(r,t)$ $:=M^{-1}\tilde{\Phi}^{\mathrm{u}}(r,t)M,\qquad r\leq t$ (23)
where $\tilde{\Phi}^{\mathrm{s}/\mathrm{u}}$ give the exponential dichotomy for (17) defined by (21).
From (23) we can directly apply Proposition 2.3 to get an exponential dichotomy $\hat{\Phi}^{\mathrm{s}/\mathrm{u}}_{*}$ for (16) in $L^{2}\!\!\times\!L^{2}$ for $r>1$ again with the same decay rate $\tilde{\eta}$, noting that the perturbed part of (16) is $\mathcal{O}(1/r)$.
Next, we use the exponential dichotomy $\hat{\Phi}^{\mathrm{s}/\mathrm{u}}_{*}$ for (16) on $L^{2}\!\!\times\!L^{2}$ and the isomorphism established by Lemma 2.4 to construct an exponential dichotomy for (9) on $X_{r}$.
Notice that the transformation $(u,v)\mapsto(A(r)^{1/2}u,v)$ and can be written as
$T(r):=\begin{pmatrix}A(r)^{1/2}&0\\ 0&1\end{pmatrix}:X_{r}\to L^{2}\!\!\times\!L^{2}\quad\text{with}\quad T(r)^{-1}=\begin{pmatrix}A(r)^{-1/2}&0\\ 0&1\end{pmatrix}:L^{2}\!\!\times\!L^{2}\to X_{r}.$
Therefore, (16) is equivalent to $\mathbf{\tilde{u}}_{r}=\big[T(r)\tilde{\mathcal{A}}(r)T(r)^{-1}+\partial_{r}\left[T(r)\right]T(r)^{-1}\big]\mathbf{\tilde{u}}$. A similar $r$-dependent conjugation by $T$ will transform $\hat{\Phi}^{\mathrm{s}}_{*}(r,t)$ and $\hat{\Phi}^{\mathrm{u}}_{*}(r,t)$ from a family of solution operators for (16) in $L^{2}\times L^{2}$ to a family of solution operators for (9) in $X^{r}$.
$\Phi^{\mathrm{s}}_{*}(r,t)$ $:=T(r)^{-1}\hat{\Phi}_{*}^{\mathrm{s}}(r,t)T(t):X_{t}\to X_{r},\qquad r\geq t$
$\Phi^{\mathrm{u}}_{*}(r,t)$ $:=T(r)^{-1}\hat{\Phi}_{*}^{\mathrm{u}}(r,t)T(t):X_{t}\to X_{r},\qquad r\leq t$ (24)
The continuity and differentiability of $\Phi^{\mathrm{s}/\mathrm{u}}_{*}$ follow directly from Lemma 2.4 and the continuity and differentiability of $\hat{\Phi}^{\mathrm{s}/\mathrm{u}}_{*}$. The projection and invariance properties required by Definition 2.2 are also inherited from $\hat{\Phi}^{\mathrm{s}/\mathrm{u}}_{*}$. All that remains to show is that $\Phi^{\mathrm{s}}_{*}$ and $\Phi^{\mathrm{u}}_{*}$ satisfy exponential decay with rate $\tilde{\eta}$; this is addressed by the second statement of Lemma 2.4.
Though the isomorphism $T(r)$ is not an isometry, by Lemma 2.4 it does respect the norm up to linear constants. Therefore the family $\Phi_{*}^{\mathrm{s}/\mathrm{u}}$ inherits the same exponential decay properties in $X_{r}$ as the exponential dichotomy $\hat{\Phi}_{*}^{\mathrm{s}/\mathrm{u}}$ has in $L^{2}\!\!\times\!L^{2}$, including the same decay rate $\tilde{\eta}$. For any $\mathbf{u}\in X_{t}$,
$\left\|\Phi_{*}^{\mathrm{s}}(r,t)\mathbf{u}\right\|_{X_{r}}^{2}$ $\leq C_{2}\left\|\hat{\Phi}_{*}^{\mathrm{s}}(r,t)T(t)\mathbf{u}\right\|_{L^{2}\!\times L^{2}}^{2}\leq C_{2}Ke^{-\tilde{\eta}(r-t)}\left\|T(t)\mathbf{u}\right\|_{L^{2}\!\times L^{2}}^{2}\leq\frac{C_{2}K}{C_{1}}e^{-\tilde{\eta}(r-t)}\left\|\mathbf{u}\right\|_{X_{t}}^{2},$ $r\geq t$
and
$\left\|\Phi_{*}^{\mathrm{u}}(r,t)\mathbf{u}\right\|_{X_{r}}^{2}$ $\leq C_{2}\left\|\hat{\Phi}_{*}^{\mathrm{u}}(r,t)T(t)\mathbf{u}\right\|_{L^{2}\!\times L^{2}}^{2}\leq C_{2}Ke^{\tilde{\eta}(r-t)}\left\|T(t)\mathbf{u}\right\|_{L^{2}\!\times L^{2}}^{2}\leq\frac{C_{2}K}{C_{1}}e^{\tilde{\eta}(r-t)}\left\|\mathbf{u}\right\|_{X_{t}}^{2},$ $r\leq t.$
Therefore (24) defines an exponential dichotomy for the principal equation (9) in $X_{r}$ for $r>1$ with decay rate $\tilde{\eta}$. Proposition 2.3 applied to (9) with perturbation $\tilde{B}(r)\in\mathcal{O}(1/r)$ from (8) gives the final result: $\Phi_{*}^{\mathrm{s}/\mathrm{u}}$ can be perturbed to an exponential dichotomy $\Phi^{\mathrm{s}/\mathrm{u}}$ for (2) with decay rate $\eta=\tilde{\eta}$. ∎
4 Conclusion.
Theorem 2.1, together with the results of [1], allow one to fully characterize the bounded behavior of spatial evolutionary systems that arise from elliptic PDEs of the form $\Delta u+V(x)u=0$ where $V(x)=V(r,\cdot)$ decays sufficiently as $r$ goes to infinity. In fact, exponential dichotomies provide a construction of bounded solutions to nonlinearly perturbed equations via a Duhamel formula. This gives rise to a constructive approach to proving the existence to localized stationary solutions of pattern-forming PDEs that lack any kind of rotational or dihedral symmetry.
To see the broad applicability of Theorem 2.1, notice that any stationary reaction diffusion model of the form $0=\Delta u+f(u)$, $u\in\mathbb{C}^{d}$, becomes $0=\Delta u+Df(q(x))u$ when linearized about a stationary solution $q(x)$. Notice also that this matrix $Df(q(x))$ will certainly decay to the constant coefficient matrix $Df(0)$ if the solution $q(x)$ is localized. Therefore the linearization of such a reaction diffusion stationary problem on $\mathbb{R}^{n}$ about any localized solution $q(x)$ has exponential dichotomies in both the core field and the far field provided that $Df(0)$ has eigenvalues away from the negative real line. We could then perturb the reaction diffusion model, including a small spatially inhomogeneous nonlinearity,
$0=\Delta u+f(u)+\mu h(x,u,\mu)=\Delta u+Df(q(x))u+\big[f(u)-Df(q(x))u+\mu h(x,u,\mu)\big]$ (25)
and use the core field and far field dichotomies for the linearized problem to construct bounded solutions to (25) in the core and far field, respectively. Showing that those two solutions connect to form a single bounded solution on the whole domain requires more work.
This type of construction was demonstrated in the channel setting, $(x,y)\in\mathbb{R}\times\Omega$, by Peterhof, Sandstede, and Scheel in [11] for a spatial evolutionary equation of the form:
$u_{x}=Au+G(u)+\mu H(x,u,\mu),\qquad u\in X=L^{2}(\Omega),$ (26)
where $A$ is an autonomous, linear (differential) operator on $L^{2}(\Omega)$ and $H$ is periodic in $x$. They assume that a homoclinic solution $q(x)$ exists for $\mu=0$ in (26), and that the linearization about $q$ has an exponential dichotomy on $\mathbb{R}^{+}$ and on $\mathbb{R}^{-}$. They proceed by constructing a bounded solution on $\mathbb{R}^{+}$ and $\mathbb{R}^{-}$ respectively and proving via a Melnikov argument that the stable space on $\mathbb{R}^{+}$ and the unstable space on $\mathbb{R}^{-}$ intersect in such a way that a unique bounded solution exists on the whole domain $\mathbb{R}$ when $\mu>0$ is small.
Ongoing and future work aims to extend this Melnikov argument to our higher dimensional context with the primary challenge being the higher dimensional kernel that comes from having not just one translational symmetry but $n$ translation symmetries, one for each unbounded direction in state space. Once such an argument is completed, the existence of exponential dichotomies in the setting described above would allow us to take previously understood radially symmetric localized solutions and perturb them in non-radially symmetric ways to prove the existence of arbitrary localized stationary patterns.
Appendix
Lemma A.1.
For all $k>0$, $\mu_{k}^{1/2}\leq(1+\mu_{k})^{1/2}$. Furthermore, $\sum_{k,j}\mu_{k}\|u_{kj}\|^{2}_{\mathbb{C}^{d}}\leq\|u\|_{H^{1}}^{2}$.
Lemma A.2.
Given (H1), for each $k>0$, $1\leq\ell\leq d$, and any arbitrary $r_{1}\geq r_{2}\geq 1$, then letting $\lambda_{\ell}={\lambda_{\ell}}^{\mathrm{re}}+\mathrm{i}{\lambda_{\ell}}^{\mathrm{im}}$,
$|\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2})|\geq C\left[({\lambda_{\ell}}^{\mathrm{re}}+\mu_{k}/{r_{2}}^{2})^{2}+\big({\lambda_{\ell}}^{\mathrm{im}}\big)^{2}\right]^{1/4}.$
Proof.
Write $\gamma_{k\ell}(r_{1})^{2}=\lambda_{\ell}+\mu_{k}/r_{1}^{2}=A_{1}e^{\mathrm{i}\theta_{1}}$ and $\gamma_{k\ell}(r_{2})^{2}=\lambda_{\ell}+\mu_{k}/r_{2}^{2}=A_{2}e^{\mathrm{i}\theta_{2}}$. Recall that $\lambda_{\ell}\in\mathbb{C}\backslash(-\infty,0]$ and therefore, since $\mu_{k}\in[0,\infty)$, $\theta_{1},\theta_{2}\in(-\pi+\varepsilon,\pi-\varepsilon)$ for some $\varepsilon>0$. Consequently, $\cos(\theta_{j}/2)\geq C>0$ for each $j=1,2$ where $C$ is some constant. We then estimate
$|\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2})|$ $=\Big|A_{1}^{1/2}\cos(\theta_{1}/2)+A_{2}^{1/2}\cos(\theta_{2}/2)+\mathrm{i}\big[A_{1}^{1/2}\sin(\theta_{1}/2)+A_{2}^{1/2}\sin(\theta_{2}/2)\big]\Big|$
$\geq A_{1}^{1/2}\cos(\theta_{1}/2)+A_{2}^{1/2}\cos(\theta_{2}/2)$
$\geq C(A_{1}^{1/2}+A_{2}^{1/2})\geq CA_{2}^{1/2}=C\left[\left({\lambda_{\ell}}^{\mathrm{re}}+\mu_{k}/{r_{2}}^{2}\right)^{2}+\big({\lambda_{\ell}}^{\mathrm{im}}\big)^{2}\right]^{1/4}$ (27)
Lemma A.3.
Suppose that either $A>0$ or $B\neq 0$, then for all $\xi\geq 0$,
$f(\xi):=\frac{\xi}{\left[\left(A+\xi^{2}\right)^{2}+B^{2}\right]^{1/4}}\leq C$
uniformly for some constant $C$.
Lemma A.4.
Given (H2), for each $k>0$, $1\leq\ell\leq d$, and $r_{1}\geq r_{2}\geq 1$ sufficiently close so that $r_{1}\leq 2r_{2}$,
$|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})|\leq C(1+\mu_{k})^{1/2}|r_{1}-r_{2}|.$
Proof.
We begin by rewriting $\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})$ as a single fraction:
$|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})|$ $=\left|\frac{\mu_{k}(r_{1}+r_{2})(r_{1}-r_{2})}{r_{1}^{2}r_{2}^{2}(\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2}))}\right|$
$\leq|r_{1}-r_{2}|\left|\frac{r_{1}+r_{2}}{r_{1}^{2}r_{2}}\right|\left|\frac{\mu_{k}}{(\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2}))}\right|.$
We next apply Lemma A.2 to bound the denominator, and apply our assumption about the proximity of $r_{1}\geq r_{2}$ to bound the coefficients via the estimates $r_{1}+r_{2}\leq 3r_{2}$ and $r_{1}^{2}r_{2}^{2}\geq r_{2}^{4}$. This yields
$|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})|$ $\leq C|r_{1}-r_{2}|\left|\frac{\mu_{k}/r_{2}^{3}}{\left[\left({\lambda_{\ell}}^{\mathrm{re}}+\mu_{k}/r_{2}^{2}\right)^{2}+\big({\lambda_{\ell}}^{\mathrm{im}}\big)^{2}\right]^{1/4}}\right|.$ (28)
By factoring $\mu_{k}/r_{2}^{3}$ into $(\mu_{k}^{1/2}/r_{2}^{2})\cdot(\mu_{k}^{1/2}/r_{2})$ and applying Lemmas A.1 and A.3, we then obtain
$|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})|$ $\leq C|r_{1}-r_{2}|\left|\mu_{k}^{1/2}/r_{2}^{2}\right|\leq C|r_{1}-r_{2}|(1+\mu_{k})^{1/2}.$
Lemma A.5.
Let $r\geq 1$ and consider $r^{\prime}\geq 1$ sufficiently close to $r$ so that $r^{\prime}/2\leq r<2r^{\prime}$. Then, given (H2), for all $k>0$ and $1\leq\ell\leq d$,
$|\gamma_{k\ell}(r)-\gamma_{k\ell}(r^{\prime})-\partial_{r}\gamma_{k\ell}(r^{\prime})(r-r^{\prime})|\leq C(1+\mu_{k})^{1/2}|r-r^{\prime}|^{2}$
where $\partial_{r}\gamma_{k\ell}$ is the $r$ derivative of $\gamma_{k\ell}$ and also the symbol for the Fréchet derivative of $A(r)^{1/2}$.
Proof.
The result follows from a rather tedious calculation in which we suppress the $\ell$ subscript:
$\gamma_{k}(r)$- $\gamma_{k}(r^{\prime})-\partial_{r}\gamma_{k}(r^{\prime})(r-r^{\prime})=\frac{\mu_{k}(1/r^{2}-1/r^{\prime 2})}{\gamma_{k}(r)+\gamma_{k}(r^{\prime})}-\frac{\mu_{k}(r^{\prime}-r)}{(r^{\prime})^{3}\gamma_{k}(r^{\prime})}$
$=(r^{\prime}-r)\frac{\mu_{k}}{(r^{\prime})^{2}}\left[\frac{r^{\prime}+r}{r^{2}(\gamma_{k}(r)+\gamma_{k}(r^{\prime}))}-\frac{1}{r^{\prime}\gamma_{k}(r^{\prime})}\right]$
$=(r^{\prime}-r)\frac{\mu_{k}}{(r^{\prime})^{2}}\left[\frac{r^{\prime}(r+r^{\prime})\gamma_{k}(r^{\prime})-r^{2}(\gamma_{k}(r)+\gamma_{k}(r^{\prime}))}{r^{\prime}r^{2}\gamma_{k}(r^{\prime})(\gamma_{k}(r^{\prime})+\gamma_{k}(r))}\right]$
$=(r^{\prime}-r)\frac{\mu_{k}}{(r^{\prime})^{2}}\left[\frac{(r^{\prime})^{2}(\gamma_{k}(r^{\prime})-\gamma_{k}(r))+\gamma_{k}(r)((r^{\prime})^{2}-r^{2})+\gamma_{k}(r^{\prime})r(r^{\prime}-r)}{r^{\prime}r^{2}\gamma_{k}(r^{\prime})(\gamma_{k}(r^{\prime})+\gamma_{k}(r))}\right]\quad\bigg(\begin{matrix}\text{add and subtract}\\ (r^{\prime})^{2}\gamma_{k}(r)\end{matrix}\bigg)$
$=-(r^{\prime}-r)^{2}\frac{\mu_{k}}{(r^{\prime})^{2}}\Bigg[\frac{(r+r^{\prime})\mu_{k}}{r^{\prime}r^{4}\gamma_{k}(r^{\prime})(\gamma_{k}(r^{\prime})+\gamma_{k}(r))^{2}}-\frac{\gamma_{k}(r)(r^{\prime}+r)}{r^{\prime}r^{2}\gamma_{k}(r^{\prime})(\gamma_{k}(r^{\prime})+\gamma_{k}(r))}$
$\qquad\qquad\qquad\qquad\quad-\frac{1}{r^{\prime}r(\gamma_{k}(r^{\prime})+\gamma_{k}(r))}\Bigg]$
$=-(r-r^{\prime})^{2}\frac{\mu_{k}^{1/2}}{r^{\prime}}\Bigg[\frac{r^{\prime}(r+r^{\prime})}{r^{4}}\frac{\mu_{k}^{3/2}/(r^{\prime})^{3}}{\gamma_{k}(r^{\prime})(\gamma_{k}(r^{\prime})+\gamma_{k}(r))^{2}}-\frac{(r+r^{\prime})}{r^{\prime}r^{2}}\frac{\gamma_{k}(r)(\mu_{k}^{1/2}/r^{\prime})}{\gamma_{k}(r^{\prime})(\gamma_{k}(r^{\prime})+\gamma_{k}(r))}$
$\qquad\qquad\qquad\qquad\quad-\frac{1}{r^{\prime}r}\frac{\mu_{k}^{1/2}/r^{\prime}}{\gamma_{k}(r^{\prime})+\gamma_{k}(r)}\Bigg]$
[…]
We can then bound $|I|$, $|II|$, and $|III|$ using the same approach as in the proof of Lemma A.4. In particular, using Lemma A.2, Lemma A.3, and the relationship between $r$ and $r^{\prime}$ we obtain:
$|I|$ $\leq 3\cdot 2^{4}\left|\frac{\mu_{k}^{3/2}/r^{\prime 2}}{\gamma_{k}(r^{\prime})(\gamma_{k}(r^{\prime})+\gamma_{k}(r))^{2}}\right|\leq C$
$|II|$ $\leq 3\cdot 4\left|\frac{\gamma_{k}(r)}{\gamma_{k}(r^{\prime})}\right|\left|\frac{\mu_{k}^{1/2}/r^{\prime}}{(\gamma_{k}(r^{\prime})+\gamma_{k}(r))}\right|\leq C$
$|III|$ $\leq 2\left|\frac{\mu_{k}^{1/2}/r^{\prime}}{(\gamma_{k}(r^{\prime})+\gamma_{k}(r))}\right|\leq C$
for some constant $C>0$ uniformly in $k$. ∎
Lemma A.6.
Given (H2), for any arbitrary $r_{1}\geq r_{2}\geq 1$ with $r_{1}$ and $r_{2}$ sufficiently close so that $r_{2}/2\leq r_{1}\leq 2r_{2}$,
$|\partial_{r}\gamma_{k\ell}(r_{1})-\partial_{r}\gamma_{k\ell}(r_{2})|\leq C(1+\mu_{k})^{1/2}\left(|r_{1}-r_{2}|+|r_{1}^{2}-r_{2}^{3}|\right).$
Proof.
We begin by writing the difference as a single fraction and simplifying.
$|\partial_{r}\gamma_{k\ell}(r_{1})-\partial_{r}\gamma_{k\ell}(r_{2})|$ $=\mu_{k}\left|\frac{1}{r_{2}^{3}\gamma_{k\ell}(r_{2})}-\frac{1}{r_{1}^{3}\gamma_{k\ell}(r_{1})}\right|=\mu_{k}\left|\frac{r_{1}^{3}\gamma_{k\ell}(r_{1})-r_{2}^{3}\gamma_{k\ell}(r_{2})}{r_{1}^{3}r_{2}^{3}\gamma_{k\ell}(r_{2})\gamma_{k\ell}(r_{1})}\right|$
$=\mu_{k}\left|\frac{r_{1}^{3}(\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2}))-\gamma_{k\ell}(r_{2})(r_{2}^{3}-r_{1}^{3})}{r_{1}^{3}r_{2}^{3}\gamma_{k\ell}(r_{2})\gamma_{k\ell}(r_{1})}\right|\qquad\text{(add and subtract $r_{1}^{3}\gamma_{k\ell}(r_{2})$ )}$
$\leq\left(\frac{\mu_{k}}{r_{2}^{3}|\gamma_{k\ell}(r_{2})\gamma_{k\ell}(r_{1})|}\right)\left|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})\right|+\left(\frac{\mu_{k}}{r_{1}^{3}r_{2}^{3}|\gamma_{k\ell}(r_{1})|}\right)\left|r_{2}^{3}-r_{1}^{3}\right|$
Applying Lemma A.2 for $r_{1}=r_{2}$, we obtain
$2|\gamma_{k\ell}(r_{j})|\geq\tilde{C}\left[({\lambda_{\ell}}^{\mathrm{re}}+\mu_{k}/r_{j}^{2})^{2}+\big({\lambda_{\ell}}^{\mathrm{im}}\big)^{2}\right]^{1/4}$
$\text{and}\qquad\frac{\mu_{k}^{1/2}/r_{j}}{|\gamma_{k\ell}(r_{j})|}\leq 2\cdot\frac{\mu_{k}^{1/2}/r_{j}}{\tilde{C}\left[({\lambda_{\ell}}^{\mathrm{re}}+\mu_{k}/r_{j}^{2})^{2}+\big({\lambda_{\ell}}^{\mathrm{im}}\big)^{2}\right]^{1/4}}$ $\leq C$
for each $j=1,2$. Furthermore, by our assumption that $r_{1}$ and $r_{2}$ are sufficiently close, we have
$\frac{\mu_{k}^{1/2}/r_{2}}{\gamma_{k\ell}(r_{1})}\leq\frac{2\mu_{k}^{1/2}/r_{1}}{\gamma_{k\ell}(r_{1})}\leq 2C.$
From the above estimates, we can further bound $\partial_{r}\gamma_{k\ell}(r_{1})-\partial_{r}\gamma_{k\ell}(r_{2})$ as follows:
$|\partial_{r}\gamma_{k\ell}(r_{1})-\partial_{r}\gamma_{k\ell}(r_{2})|$ $\leq C\left[\left(\frac{\mu_{k}^{1/2}}{r_{2}^{2}|\gamma_{k\ell}(r_{1})|}\right)\left|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})\right|+\left(\frac{\mu_{k}^{1/2}}{r_{1}^{2}r_{2}^{3}}\right)\left|r_{1}^{3}-r_{2}^{3}\right|\right]$
$\leq C\left[\left(\frac{2C}{r_{2}}\right)\left|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})\right|+\left(\frac{\mu_{k}^{1/2}}{r_{1}^{2}r_{2}^{3}}\right)\left|r_{1}^{3}-r_{2}^{3}\right|\right]$
$\leq C\left[2C\left|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})\right|+\mu_{k}^{1/2}\left|r_{1}^{3}-r_{2}^{3}\right|\right],\qquad\quad(r_{1},r_{2}\geq 1).$
Lastly, we apply Lemmas A.1 and A.4 to obtain:
$|\partial_{r}\gamma_{k\ell}(r_{1})-\partial_{r}\gamma_{k\ell}(r_{2})|$ $\leq C(1+\mu_{k})^{1/2}\left(|r_{1}-r_{2}|+|r_{1}^{3}-r_{2}^{3}|\right)$
Lemma A.7.
For any $u\in H^{1}(S^{n-1},\mathbb{C})$, there is a constant $C>0$ and independent of $r$ such that
$\frac{1}{r^{2}}\|u\|_{H^{1}}^{2}\leq C\|u\|_{L^{2}}^{2}+\|A(r)^{1/2}u\|_{L^{2}}^{2}.$
Furthermore, if (H2) holds, then for some constant $C$ independent of $r$,
$\|u\|_{L^{2}}^{2}\leq C\|A(r)^{1/2}u\|_{L^{2}}^{2}.$
Proof.
By expanding the $H^{1}$ norm of $u$, adding and subtracting $r^{2}V_{\infty}$, and applying triangle inequality we obtain
$\frac{1}{r^{2}}\left\|u\right\|_{H^{1}}^{2}=\sum_{k,j}\bigg\|\bigg(\frac{1+\mu_{k}}{r^{2}}\bigg)u_{kj}\bigg\|^{2}_{\mathbb{C}^{d}}$ $\leq\sum_{k,j}\left\|\left(\frac{1-r^{2}V_{\infty}}{r^{2}}+\frac{r^{2}V_{\infty}+\mu_{k}}{r^{2}}\right)u_{kj}\right\|_{\mathbb{C}^{d}}^{2}$
$\leq\sum_{k,j}\left(1+\|V_{\infty}\|\right)\left\|u_{kj}\right\|^{2}_{\mathbb{C}^{d}}+\sum_{k,j}\left\|A_{k}(r)^{1/2}u_{kj}\right\|^{2}_{\mathbb{C}^{d}}$
$=C\left\|u\right\|_{L^{2}}^{2}+\left\|A(r)^{1/2}u\right\|_{L^{2}}^{2}.$
If (H2) holds, then we have an $r$-independent lower bound on $|\gamma_{k\ell}(r)|^{2}$ given by two cases:
$|\gamma_{k\ell}(r)|^{2}=|\gamma_{k\ell}(r)^{2}|=\sqrt{({\lambda_{\ell}}^{\mathrm{re}}+\mu_{k}/r^{2})^{2}+\big({\lambda_{\ell}}^{\mathrm{im}}\big)^{2}}\geq\begin{cases}\left|{\lambda_{\ell}}^{\mathrm{im}}\right|&\text{ if }{\lambda_{\ell}}^{\mathrm{im}}\neq 0\\ \,\left|{\lambda_{\ell}}^{\mathrm{re}}\right|&\text{ if }{\lambda_{\ell}}^{\mathrm{im}}=0\end{cases}$ (29)
where we note that the inequality $|\gamma_{k\ell}(r)|^{2}\geq|\lambda_{\ell}^{\mathrm{re}}|$ holds by (H2) only in the case where $\lambda_{\ell}^{\mathrm{im}}=0$.
If $d=1$, then (29) directly implies $\|A(r)^{1/2}u\|_{L^{2}}^{2}\geq C\|u\|_{L^{2}}^{2}$ for a constant $C>0$ independent of $r$.
If $d>1$, then (29) means that $|\gamma_{k\ell}(r)|$ is bounded away from zero uniformly in $k$, $\ell$, and $r$. Let $\Gamma>0$ be this uniform lower bound for all $|\gamma_{k\ell}(r)|$. We can now use a similar argument as that used in Equation (20) to bound $\|u\|_{L^{2}}$ in terms of these scalar coefficients. To do this, recall that $R$ is the matrix whose columns are the eigenvector basis $\{\rho_{1},\ldots,\rho_{d}\}$ for $\mathbb{C}^{d}$ associated with $V_{\infty}$. We derive the desired bound by rewriting $u$ in terms of the eigenvector basis as follows:
$\big\|u\big\|_{L^{2}}^{2}$ $=\sum_{k,j}\|RR^{-1}u_{kj}\|_{\mathbb{C}^{d}}^{2}\leq\sum_{k,j}\|R\|^{2}\big\|R^{-1}u_{kj}\big\|_{\mathbb{C}^{d}}^{2}$
$=\sum_{k,j}\|R\|^{2}\sum_{\ell=1}^{d}|u_{kj\ell}|^{2}\leq\sum_{k,j}\|R\|^{2}\sum_{\ell=1}^{d}\frac{|\gamma_{k\ell}(r)|^{2}}{\Gamma^{2}}|u_{kj\ell}|^{2}$
$=\sum_{k,j}\frac{\|R\|^{2}}{\Gamma^{2}}\left\|\begin{pmatrix}\gamma_{k1}(r)&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\gamma_{kd}(r)\end{pmatrix}R^{-1}u_{kj}\right\|^{2}_{\mathbb{C}^{d}}$
$=\sum_{k,j}\frac{\|R\|^{2}}{\Gamma^{2}}\left\|R^{-1}A_{k}(r)^{1/2}u_{kj}\right\|^{2}_{\mathbb{C}^{d}}\leq\frac{\|R\|^{2}\|R^{-1}\|^{2}}{\Gamma^{2}}\left\|A(r)^{1/2}u\right\|_{\mathbb{C}^{d}}^{2}.$
Lemma A.8.
The real part of $\gamma_{k\ell}(r)$ is bounded below by $\mathrm{Re}\left[\lambda_{\ell}^{1/2}\right]$.
$\mathrm{Re}\!\left[\gamma_{k\ell}(r)\right]=\mathrm{Re}\!\left[\left(\lambda_{\ell}+\mu_{k}/r^{2}\right)^{1/2}\right]\geq\mathrm{Re}\!\left[\lambda_{\ell}^{1/2}\right]$
Proof.
If $\lambda_{\ell}\in\mathbb{R}$, then $\lambda_{\ell}>0$ by hypothesis (H2) and the result follows directly.
If $\lambda_{\ell}^{\mathrm{im}}\neq 0$, denote the argument of $\gamma_{k\ell}(r)^{2}$ by $\theta_{k\ell}(r)\in[-\pi,\pi)$ so that $\gamma_{k\ell}(r)^{2}=|\gamma_{k\ell}(r)^{2}|e^{i\theta_{k\ell}(r)}$. Similarly, let $\theta_{\ell}$ be the argument of $\lambda_{\ell}$. For the remainder of this proof, assume without loss of generality that $\lambda_{\ell}^{\mathrm{im}}>0$ so that $\theta_{\ell}\in(0,\pi)$. Since $\mu_{k}\geq 0$ and $\gamma_{k\ell}(r)^{2}=\lambda_{\ell}+\mu_{k}/r^{2}$, we have that $\theta_{k\ell}(r)\in(0,\theta_{\ell}]$. Furthermore, since $\mathrm{Im}\left[\gamma_{k\ell}(r)^{2}\right]=\lambda_{\ell}^{\mathrm{im}}$ for all $k$ and $r$, we have that $\left|\gamma_{k\ell}(r)^{2}\right|=\frac{\lambda_{\ell}^{\mathrm{im}}}{\sin(\theta_{k\ell}(r))}$ and we can write the real part of $\gamma_{k\ell}(r)$ directly as a function $A$ of $\theta_{k\ell}(r)$:
$\mathrm{Re}\left[\gamma_{k\ell}(r)\right]=\sqrt{\lambda_{\ell}^{\mathrm{im}}}A(\theta_{k\ell}(r)),\qquad A(\theta):=\frac{\cos(\theta/2)}{\sqrt{\sin(\theta)}},\,\,\theta\in(0,\pi)$
Notice that $\mathrm{Re}[\lambda_{\ell}^{1/2}]=\sqrt{\lambda_{\ell}^{\mathrm{im}}}A(\theta_{\ell})$, and recall that $\theta_{k\ell}(r)\in(0,\theta_{\ell})$. To show that $\mathrm{Re}\left[\gamma_{k\ell}(r)\right]\geq\mathrm{Re}[\lambda_{\ell}^{1/2}]$, it is therefore sufficient to show that the function $A(\theta)$ is decreasing on $(0,\theta_{\ell}]$.
Taking a derivative, applying product rule and simplifying with double-angle trigonometric identities,
$A^{\prime}(\theta)$ $=\frac{-\cos(\theta/2)\cos(\theta)}{2\sin^{3/2}(\theta)}+\frac{-\sin(\theta/2)}{2\sin^{1/2}(\theta)}=\frac{-\cos(\theta/2)\left[\cos(\theta)+2\sin^{2}(\theta/2)\right]}{2\sin^{3/2}(\theta)}=\frac{-\cos(\theta/2)}{2\sin^{3/2}(\theta)}$
Since both $\cos(\theta/2)$ and $\sin^{3/2}(\theta)$ are strictly positive for $\theta\in(0,\pi)$, $A^{\prime}(\theta)$ is strictly negative and $A$ is a decreasing function on $\theta\in(0,\pi)$. It follows that
$\mathrm{Re}\left[\gamma_{k\ell}(r)\right]=\sqrt{\lambda_{\ell}^{\mathrm{im}}}A(\theta_{k\ell}(r))\geq\sqrt{\lambda_{\ell}^{\mathrm{im}}}A(\theta_{\ell})=\mathrm{Re}\left[\lambda_{\ell}^{1/2}\right]\quad\text{ for all }k>0\text{ and }r>1.$